# Practicing Stress Relief for the Everyday: Designing Social Simulation Using VR, AR, and LLMs

### v1

@@ 2.3.2. Large Language Models (LLMs) for Mental Health
There has been remarkable success and growth with applying large language models to the mental health domain whether in the clinical setting or non-professional context. LLMs have been used to directly provide therapeutic help (Lai et al., 2023; Loh and Raamkumar, 2023; Fu et al., 2023), classify mental health experiences (Xu et al., 2024), and teach social and coping skills (Shaikh et al., 2024; Yang et al., 2024b; Hu et al., 2024). There is also emerging work on the capabilities of LLMs to help people gain mental health insights for themselves. For example, LLMs can help patients with their journaling content and then summarize patients’ thoughts for their caretakers (Kim et al., 2024) and help suggest restructuring of negative thoughts (Sharma et al., 2024).
@@ 3.3.1. Participants and Recruitment
We showed the eight prototypes on the participant’s highest stress topic given their initial survey. In addition, we also showed the ”onboarding” guide to the breathing guidance system before the participant saw VR-Guided or AR-Guided to help users engage with the exercise (Patibanda et al., 2017); in the onboarding guide, the participant saw only the breathing orb and no other simulation (see Figure 3), during which the participant could take as long as they needed to practice the box breathing pattern before proceeding. For each prototype, we provided an introduction blurb for participants to imagine. For example, before showing participants the VR prototypes for interpersonal conflict with a roommate, we stated: ”Imagine that you recently moved in with a new roommate. Things have been tense, though, and you plan to confront your roommate at home later today after work about an ongoing source of disagreement between you two. Nervous about how this conversation will go, you put on a virtual reality simulation to practice what you are going to say to them. In this scene, you’ll be met with your roommate who will speak with you about this ongoing conflict. You’ll be able to choose the topic of the conflict.” For the public speaking and interpersonal conflict scenes, participants could choose their own topic of choice for their speaking topic and conflict reason, respectively, or choose from pre-sets as shown in Figure 4.
@@ 3.4. Data Analysis
In addition to how participants thought of the simulation prototypes in general, we also wanted to measure participants’ preferences among the different prototype options (e.g. VR versus AR versus text-based modalities). To produce numerical ratings for the eight prototype concepts, the first two authors independently read participants’ reactions to each design concept and rated it from 1 (disliked a lot) to 5 (liked a lot); see Figure 6 for results. Note that these ratings were given by the researchers from analyzing the participants’ interviews, rather than the participants directly giving numerical ratings of the prototypes themselves. The authors had high internal consistency with a Cronbach’s alpha of 0.88, suggesting that the authors’ ratings reliably indicate participants’ preferences among the prototypes.
@@ 4.3. Modality: Benefits and Potential Risks of Immersion and Realism
In terms of which modalities were preferred among the eight prototypes, participants generally preferred AR over VR or text-based prototypes, with AR-Interactive and AR-Guided receiving the highest average ratings (3.74 out of 5) (see Figure 6). However, rather than a strict favorite, some saw value in multiple designs depending on context:
@@
One notable consensus among participants regardless of their overall preferences was that AR was most useful when available in the actual locations where they anticipated stressors, such as the classroom they expected to give a talk in or their apartment if expecting a confrontation to occur in their home. For example, participants said that they’d ”use AR when it is close to where I will actually give a speech” (P19) and practice stress-relief for a roommate conflict ”in my actual room…maybe I can breathe and walk around while I breathe” (P9). Indeed, ”a lot of the confidence comes from just being comfortable in your space” (P12) and practicing situations in the expected environment was already what some participants said they would do now if preparing for something like public speaking. This would also allow for people to practice their own physicality. For example, P9 felt AR would ”encourage me to move around a little bit, do hand gestures” and, more than the dialogue, P17 saw the simulation as more useful for his ”physicality, you know moving around…engaging with people, which approaches I would do. I could train that”.
@@ 5.1. Designing Simulation for Everyday Self-Care
Modality. One of the most significant findings was the clear preference for AR among our prototypes. Participants envisioned using AR in real-world environments to overlay both practice scenarios and mental health guidance, allowing for the direct application of coping strategies in the settings they encounter in daily life. In the context of everyday stress, where individuals are frequently exposed to environments that trigger emotional or mental distress, our interview findings suggest that AR lowers barriers in transferring skills to real-world settings and individuals prefer to practice coping techniques in spaces that closely resemble real-world settings. While high levels of immersion in VR can foster engagement and be beneficial for simulating scenarios that aren’t immediately accessible, its separation from real-world contexts may limit the practical application of skills learned in the virtual space. Text-based prototypes, unsurprisingly, were considered the least immersive given absence of visual and auditory feedback. However, many participants said they would use text-based systems because of its availability throughout the day, suggesting that even non-immersive tools have a valuable role in the self-care space.
@@
Participants’ reactions to our prototypes echoed past findings that LLMs are capable of generating natural, context-aware dialogue that participants found realistic compared to real-world interactions. In the future, given the success of speech-to-text in our study, LLMs could be used to also dynamically design environments rather than just provide dialogue as it did in our study. Given the accuracy of our speech-to-text engine using WhisperAI, it is plausible that users could verbally describe anticipated situations, and the LLM could then generate corresponding dialogue, character personas, physical avatar appearances, and environmental features to create a VR or AR scene for the user to practice with. Since participants consistently expressed a desire for customization features to adjust simulations to reflect their personal environments and stressors, using LLMs for designing more aspects of the simulation could facilitate this capability while reducing the time, effort, and technical skill needed on the user’s part.
@@ 5.3.2. Reframing HCI for Self-Care
Lastly, we discuss how we hope our work contributes to the broader landscape of HCI for mental health by framing self-care — and mental healthcare in general — as fundamental needs for everyone, not just those with diagnosable, chronic conditions. Traditionally, self-care technologies in HCI have been positioned within the framework of clinical interventions or therapeutic support, often focusing on structured tools for chronic conditions like depression or anxiety . However, as our findings highlight, there is a growing need to reframe self-care in HCI to create tools for everyday practice; in particular, we aimed to use HCI methods to equip users with the skills to manage stress in real-time for the contexts where it occurs. Our work’s findings and implications also emphasize the idea of people’s own self-efficacy in taking care of themselves. As seen in our interviews, people want to be self-sufficient but do not currently engage in means to do so when it comes to mental health, despite it being crucial for our well-being just as much (and maybe even more) as social or professional support needs. In our study, participants were able to engage in simulation that they felt gave agency to learning self-care skills in the long-term to enact on their own in the real-world.
In addition, self-care in HCI has traditionally focused on reactive strategies to help people manage stress after it occurs or self-management and tracking for specific conditions and diseases . However, our study suggests that HCI research can go beyond these framings to instead foster proactive self-care by allowing users to practice personal coping skills before they enter stressful situations. This shifts self-care from a purely reactive measure to an ongoing skill-building process, where users can refine their emotional regulation techniques over time. HCI’s role in supporting this proactive approach is critical, as it positions technology not just as a fix but as an integrated part of daily life that empowers users to actively manage their mental health on their own terms.

### v3

@@ 2.3.2. Large Language Models (LLMs) for Mental Health
There has been remarkable success and growth with applying large language models to the mental health domain whether in the clinical setting or non-professional context. LLMs have been used to directly provide therapeutic help (Lai et al., 2023; Loh and Raamkumar, 2023; Fu et al., 2023), classify mental health experiences (Xu et al., 2024), and teach social and coping skills (Shaikh et al., 2024; Yang et al., 2024b; Hu et al., 2024). There is also emerging work on the capabilities of LLMs to help people gain mental health insights for themselves. For example, LLMs can help patients with their journaling content and then summarize patients’ thoughts for their caretakers (Kim et al., 2024a) and help suggest restructuring of negative thoughts (Sharma et al., 2024).
@@ 3.3.1. Participants and Recruitment
We showed the eight prototypes on the participant’s highest stress topic given their initial survey, in the progressive order listed in Section 3.2 (VR-Static, then VR-Interactive, then VR-Guided, …). We followed this order so that participants had a progressive familiarity and learning of additional features, from no interaction, to added interaction, and finally added guidance features. In addition, we also showed the ”onboarding” guide to the breathing guidance system before the participant saw VR-Guided or AR-Guided to help users engage with the exercise (Patibanda et al., 2017); in the onboarding guide, the participant saw only the breathing orb and no other simulation (see Figure 3), during which the participant could take as long as they needed to practice the box breathing pattern before proceeding. For each prototype, we provided an introduction blurb for participants to imagine. For example, before showing participants the VR prototypes for interpersonal conflict with a roommate, we stated: ”Imagine that you recently moved in with a new roommate. Things have been tense, though, and you plan to confront your roommate at home later today after work about an ongoing source of disagreement between you two. Nervous about how this conversation will go, you put on a virtual reality simulation to practice what you are going to say to them. In this scene, you’ll be met with your roommate who will speak with you about this ongoing conflict. You’ll be able to choose the topic of the conflict.” For the public speaking and interpersonal conflict scenes, participants could choose their own topic of choice for their speaking topic and conflict reason, respectively, or choose from pre-sets as shown in Figure 4.
@@ 4.3. Modality: Benefits and Potential Risks of Immersion and Realism
Rather than expressing a universal favorite among the modalities shown in the study, many participants saw value in multiple designs depending on context:
@@
One notable consensus among participants, though, was that they were most likely to use AR if they had access to the real physical locations where anticipated stressors might occur; for example, participants said that they’d ”use AR when it is close to where I will actually give a speech” (P19) such as in a classroom, and practice stress-relief for a roommate conflict ”in my actual room…maybe I can breathe and walk around while I breathe” (P9). Indeed, ”a lot of the confidence comes from just being comfortable in your space” (P12) and practicing situations in the expected environment was already what some participants said they would do now if preparing for something like public speaking. This would also allow for people to practice their own physicality. For example, P9 felt AR would ”encourage me to move around a little bit, do hand gestures” and, more than the dialogue, P17 saw the simulation as more useful for his ”physicality, you know moving around…engaging with people, which approaches I would do. I could train that”.
@@ 5.1. Designing Simulation for Everyday Self-Care
Modality. One of the most significant findings was the clear preference for AR among our prototypes. Participants envisioned using AR in real-world environments to overlay both practice scenarios and mental health guidance, allowing for the direct application of coping strategies in the settings they encounter in daily life. In the context of everyday stress, where individuals are frequently exposed to environments that trigger emotional or mental distress, our interview findings suggest that AR lowers barriers in transferring skills to real-world settings and individuals prefer to practice coping techniques in spaces that closely resemble real-world settings. While high levels of immersion in VR/AR can foster engagement and be beneficial for simulating scenarios that aren’t immediately accessible (Shin, 2019; Bowman and McMahan, 2007), its separation from real-world contexts may limit the practical application of skills learned in the virtual space. Text-based prototypes, unsurprisingly, were considered the least immersive given absence of visual and auditory feedback. However, many participants said they would use text-based systems because of its availability throughout the day, suggesting that even non-immersive tools have a valuable role in the self-care space.
@@
Participants’ reactions to our prototypes echoed past findings that LLMs are capable of generating natural, realistic dialogue (Gray et al., 2024; Kim et al., 2024b; Duan et al., 2023) that participants found comparable to their expectations of real-world interactions. In the future, given the success of speech-to-text in our study, LLMs could be used to also dynamically design environments rather than just provide dialogue as it did in our study. Given the accuracy of our speech-to-text engine using WhisperAI, it is plausible that users could verbally describe anticipated situations, and the LLM could then generate corresponding dialogue, character personas, physical avatar appearances, and environmental features to create a VR or AR scene for the user to practice with. Since participants consistently expressed a desire for customization features to adjust simulations to reflect their personal environments and stressors, using LLMs for designing more aspects of the simulation could facilitate this capability while reducing the time, effort, and technical skill needed on the user’s part.
@@ 5.3.2. Reframing HCI for Self-Care
Lastly, we discuss how we hope our work contributes to the broader landscape of HCI for mental health by framing self-care — and mental healthcare in general — as fundamental needs for everyone, not just those with diagnosable, chronic conditions. Traditionally, self-care technologies in HCI have been positioned within the framework of clinical interventions or therapeutic support, often focusing on structured tools for chronic conditions like depression or anxiety (Nunes et al., 2015). However, as our findings highlight, there is a growing need to reframe self-care in HCI to create tools for everyday practice; in particular, we aimed to use HCI methods to equip users with the skills to manage stress in real-time for the contexts where it occurs. Our work’s findings and implications also emphasize the idea of people’s own self-efficacy in taking care of themselves. As seen in our interviews, people want to be self-sufficient but do not currently engage in means to do so when it comes to mental health, despite it being crucial for our well-being just as much (and maybe even more) as social or professional support needs. In our study, participants were able to engage in simulation that they felt gave agency to learning self-care skills in the long-term to enact on their own in the real-world.
Self-care in HCI has traditionally focused on strategies including bringing detection and awareness to stress (Neigel et al., 2024; Giannakakis et al., 2019; Akiri et al., 2024; Panicker and Gayathri, 2019) or tracking symptoms for specific conditions and diseases (Nunes et al., 2015; Wang et al., 2018; Beltzer et al., 2023). However, our study suggests that HCI research can go beyond these framings to instead foster proactive self-care by allowing users to practice personal coping skills for applying during future situations. This shifts self-care from a purely awareness or reactive measure to an ongoing skill-building process, where users can refine their emotional regulation techniques over time. HCI’s role in supporting this proactive approach is critical, as it positions technology not just as a fix but as an integrated part of daily life that empowers users to actively manage their mental health on their own terms.
6. Limitations and Future Work
Order effects – including recency affects and fatigue, which could be of notable concern particularly in studies using VR/AR technology which can easily fatigue participants.